# countnullcountdimennullnullnulldimennullnullnulldimennullnullnulldimennullnullnulldimen

### v1

@@ 1.1 Our contributions
  • Our result creates a surprising contrast with the mean-variance ambiguity set, where a straightforward concentration argument yields a common value of $\mu$.
  • For deterministic mechanisms and fixed $m$, the maximin and minimax problem typically do not have the same value; compare, e.g., [42, 24] for the case $m=1$. One has to resort to randomization to establish such results. [^2] For the case $m=1$, the fact that the maximin and minimax problem have the same value when randomization is allowed, follows from [19, Lemma 1]. To the best of our knowledge no such statement is known for the independent multi-good setting. As $m$ grows large, we show that the value gap disappears for deterministic mechanisms and the problems are equivalent in terms of revenue.
  • Solving the maximin and minimax problem under mean-MAD ambiguity with independent good distributions is arguably much harder than its correlated counterpart, where one can resort to well-known primal-dual semi-infinite linear programming techniques or computing a saddle point of a zero-sum game. Computing optimal independent distributions, from an optimization point of view, results in non-linear (and non-convex) problems. A fitting analogy is comparing the difference in computational complexity between computing a mixed Nash equilibrium (difficult) and a correlated equilibrium in general finite games (tractable via a linear program).
In our technical analysis the subset $\mathcal{P}_{2}(\mu,d)\subseteq\mathcal{P}(\mu,d)$ of all distributions supported on two points plays an important role. In fact, Theorem 1.1 is also true for this ambiguity set:
Corollary 1.2.
The statement of Theorem 1.1 remains true if we replace $\mathcal{P}(\mu,d)$ by $\mathcal{P}_{2}(\mu,d)$: The set of all non-negative distributions with mean $\mu$ and MAD $d$ whose support consists of two points.
For the case $\mu\leq d<2\mu$, which turns out to be more complex, we have the partial result showing that the bundling mechanism achieves a value of $\mu-d/2$ and that no deterministic mechanism can do better than $d/2$. We use $\mathrm{BUND}(p_{m},\mathbb{P}^{m})$ to denote the revenue of the bundling mechanism that sets price $p_{m}$ for selling $m$ independent goods with joint distribution $\mathbb{P}^{m}$ (see Section 2.2 for the formal definition). The first and second inequality in Theorem 1.3 trivially hold.
Theorem 1.3.
For $0<\mu\leq d<2\mu$, it holds that
@@
$\mu-d/2$ $=\lim_{m\rightarrow\infty}\displaystyle\sup_{p_{m}\geq 0}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\leq\lim_{m\rightarrow\infty}\,\displaystyle\sup_{D\in\mathcal{D}_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}$
$\displaystyle\leq\lim_{m\rightarrow\infty}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}\,}\,\sup_{D\in\mathcal{D}_{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}\leq d/2.$ (1.5)
To prove the above theorems, we first provide results for the maximin and minimax problem separately in Section 3 and conclude with a summary of how Theorems 1.1 and 1.3 follow from them. We will show that for any $d$ the bundling mechanism yields a revenue of $\mu-d/2$ in the maximin problem and that no algorithm can guarantee more than $\max\{\mu-d/2,d/2\}$ for the minimax problem.
One question that arises is what the true answer is for the range $\mu\leq d<2\mu$. Is there a gap between the revenue guarantee of the maximin and the minimax problem? We show in Proposition E.1 in Appendix E that for the set of two-point distributions $\mathcal{P}_{2}(\mu,d)$ and i.i.d. goods, the minimax problem for the bundling mechanism gives a value strictly greater than $\mu-d/2$, which indicates that the second part of Theorem 1.1 (i.e., Corollary 1.2) is not true for the range $\mu\leq d<2\mu$. We therefore conjecture that $\mu-d/2$ is not the true value of the minimax problem for $\mu\leq d<2\mu$ for the ambiguity set $\mathcal{P}(\mu,d)$. All together, our results and proofs show an intriguing dependence on the dispersion parameter $d$ that deserves further investigation.
Inspired by the optimality of the bundling mechanism, we further investigate this mechanism for two other revenue objectives from the literature: the ratio objective (e.g., [28]) and the absolute regret objective (e.g, [18]), defined, respectively, by
$\lim_{m\rightarrow\infty}\sup_{p_{m}}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}$ (1.6)
and
$\lim_{m\rightarrow\infty}\inf_{p_{m}}\,\sup_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}.$ (1.7)
Here $\text{OPT}(m,\mathbb{P}^{m})$ is the expected revenue of the (unknown) optimal deterministic mechanism for selling $m$ independent goods with known joint distribution $\mathbb{P}^{m}$. When $m=1$, the absolute revenue, ratio, and absolute regret objectives all have a different optimal posted price for the problems in (1.2), (1.6), and (1.7) under mean-MAD ambiguity [42, 31, 45]. We show that, when $m\rightarrow\infty$, a bundling price $p_{m}\approx m(\mu-d/2)$ is optimal for both the ratio (Theorem 4.1) and absolute regret objective (Theorem 4.2) among all bundling prices, just as it was for the absolute revenue objective (Theorems 1.1 and 1.3). This shows that all three objectives that we consider in fact have the same optimal bundling price as $m$ grows large, which stands in stark contrast with the case $m=1$. Roughly speaking, our results give rise to a universality of the bundling price $p_{m}\approx m(\mu-d/2)$ as an optimal robust solution among various revenue objectives.
An interesting direction for future research is to understand how allowing randomization in the seller’s mechanism affects our results. In particular, can one improve over the value of $\mu-d/2$ for the absolute revenue objective, and the corresponding values of the other objectives? Secondly, (dis)proving that the bundling mechanism is optimal among all deterministic mechanisms for the ratio and absolute regret objective is also a very intriguing question.
Techniques. One of the main technical tools that we use to establish our results is a one-sided concentration inequality for distributions with known mean and mean absolute deviation (Proposition 3.1), which might be of independent interest. To establish our tight result in Theorem 1.1 we lower bound the maximin problem with a value of $\mu-d/2$, by analysing the bundling mechanism, and upper bound the minimax problem by $\max\{d/2,\mu-d/2\}$. The latter is the technically most challenging part. We establish this bound by analysing the minimax problem for two-point distributions, and argue that in this case only the smallest two support points of $\sum_{i}X_{i}$, the sum of the valuations of the buyer, are relevant for the analysis. Using the definition of a truthful mechanism, we then argue that an optimal mechanism is one that chooses a bundling price to be the better choice of those two support points (the first giving a revenue of $\mu-d/2$ and the second one $d/2$).
1.2 Further related work
In this section we discuss related works, starting with the classical Bayesian setting. We then continue with robust mechanisms for selling multiple goods in other settings, as well as the literature that addresses the case when there are multiple buyers. Finally, we discuss related works for robustly selling a single good. These works demonstrate the growing interest in obtaining distributionally robust solutions to classical mechanism design problems, to which our work contributes.
Mechanisms for selling multiple goods in the classical Bayesian setting. Whereas the optimal mechanism for selling a single good is known, the multi-good setting is more complex. Natural choices like bundling all goods together or selling them separately are in general not optimal for any distribution. Furthermore, the optimal mechanism need not be deterministic, as opposed to the single-good case. These observations are nicely summarized by Hart and Nisan [30, Examples 1-4]. Daskalakis, Deckelbaum and Tzamos [22] provide a primal-dual framework that allows them to characterize the optimal mechanism for selling a fixed number of $m\geq 2$ goods, thereby also generalizing earlier works such as [34, 11, 28]; see also [12] for a more general framework. Daskalakis et al. [22] also provide a characterization for when offering the grand bundle is optimal. As a concrete application of this characterization, they consider i.i.d. goods with uniform distribution on $[c,c+1]$ for some scalar $c$. They show that if $m$ is kept fixed, then for $c$ large enough the bundling mechanism is optimal (this was shown by Pavlov [39] for $m=2$). However, for $c$ fixed, it is shown that for $m$ large enough, the bundling mechanism is not optimal for selling $m$ i.i.d. goods. [^3] We suspect that in our minimax setting bundling is also not optimal for fixed $m$, but leave this for future work. In general, there is not a simple economic interpretation of the characterization in [22]. There have also been other works concerned with deriving conditions on the known independently distributed goods for when bundling is the optimal mechanism, such as [29, 39, 22, 35, 34]. See [16, Footnote 1] for an overview.
Robust mechanisms for selling multiple goods. Caroll [14] introduces a robust perspective for selling multiple goods. It is shown that when correlations between the goods are allowed, the optimal robust mechanism is to sell every good separately. See also [15] for an interesting survey. Che and Zhong [16] generalize Caroll’s work by introducing a partition structure. They then assume to know the mean of every good, and some dispersion information about every bundle in the partition, but not the possible correlations between the goods in different parts of the partition. Deb and Roesler [23] consider a robust setting related to ours, but instead of moment information assume to have only support information for the unknown distributions, i.e., they only assume that the distributions have positive probability mass on a given interval. They prove that randomized bundling is the optimal mechanism for selling $m$ goods under an exchangeability assumption (that also applies to independent distributions); see also Che and Zhong [16] for related results in this direction. Giannakopoulos, Pocas and Tsigonias-Dimitriadis [28] consider the robust ratio objective and provide various tight, up to constants, results for selling multiple goods under mean-variance ambiguity for both correlated and independent value distributions. See also the work of Chen, Hu and Perakis [18] for results in this direction.
Robust mechanisms for multiple buyers. There are also various papers that study the setting with multiple buyers where there are unknown distributions coming from (moment-based) ambiguity sets. Anunrojwong, Balseiro and Besbes [3] study the problem of selling one good to $n$ potential buyers, where the unknown (joint) distribution is only assumed to have bounded support, meaning an upper bound on the valuations is known. Bachrach and Talgam-Cohen [5] consider a similar model for selling a single good to $n$ i.d.d. (potential) buyers, but in addition assume to know the mean of the common distribution. In both these works (in [5] only for $n=2$), a robust version of the second price auction with (randomized) reserve price arises as the optimal robust solution, which is proved by using a saddle-point argument for the corresponding robust zero-sum game between the seller and nature. For other works in the multi-buyer setting, see, e.g., Kocyigit et al. [32], Allouah and Besbes [2], or the survey of Caroll [15].
Robust mechanisms for single good. Maximin analysis for mean-variance ambiguity was pioneered in Azar and Micali [4], generalized to higher moments in [13]; see also [18, 44]. Other forms of knowledge were also studied, such as percentiles [25], mean absolute deviation [42, 24, 19], or knowing that the valuation distribution is within the proximity of a given reference distribution [7, 19]. Instead of maximin expected revenue, alternative objectives studied in the literature include minimax regret [8, 7, 24, 19] and the competitive or approximation ratio [25, 28, 24]. We also refer to the work of Allouah, Bahamou and Besbes [1] who assume to have prior knowledge about the probability of sale of a good. In some works in the literature, randomized algorithms are also considered and shown to be optimal in certain cases. For selling one good in the robust setting, Chen, Hu and Wang [19] show, based on a functional version of Von Neumann’s minimax theorem [10], that if randomized mechanisms are allowed in the maximin problem, the values of the maximin and minimax problems are equal to each other. We are not aware of a similar result for the problem of selling multiple goods, in particular not with the additional assumption of independent goods. This relation fails to hold if only (simple) deterministic mechanisms, like posted prices, are allowed.
1.3 Outline
We give formal definitions of all the relevant notions in Section 2. In Section 3 we give the proofs of Theorems 1.1 and 1.3. In Section C.2 we show that $p_{m}\approx m(\mu-d/2)$ is in fact the optimal bundling price for the ratio (Theorem 4.1) and absolute regret (Theorem 4.2) objective.
2 Preliminaries
We will start by formally defining deterministic truthful mechanisms and the bundling mechanism. We then introduce the various robust revenue objectives using probability theory.
2.1 Mechanism design
A direct revelation, deterministic mechanism $D$ in the setting of a single buyer that bids on $m$ goods is defined by a pair $(z,\pi)$. Here $z:\mathbb{R}_{\geq 0}^{m}\rightarrow\{0,1\}^{m}$ is the allocation rule, and $\pi:\mathbb{R}_{\geq 0}^{m}\rightarrow\mathbb{R}_{\geq 0}$ the payment rule. For a given valuation vector $v=(v_{1},\dots,v_{m})$ with $v_{i}$ the value the buyer has for good $i$, they receive the goods $i$ for which $z_{i}(v)=1$, and get charged $\pi(v)$ in total for all the goods they receive. The utility of the buyer under mechanism $D$ with valuation vector $v$ is then given by $u(D,v)=\langle z(v),v\rangle-\pi(v)$ with $\langle x,y\rangle=\sum_{i}x_{i}y_{i}$ the inner product for two vectors $x,y\in\mathbb{R}^{m}$.
A mechanism $D=(z,\pi)$ is truthful if the following conditions are satisfied:
The first condition ensures that the buyer has no incentive to misreport the true values $v$ that they have for the goods; that is, bidding truthfully is a dominant strategy that maximizes their utility. We refer to this property as incentive compatibility. The second condition guarantees that if a buyer truthfully reports their values, then their utility (or surplus) is non-negative, i.e., they don’t lose anything by participating in the selling mechanism. This property is called individual rationality. We use $\mathcal{D}_{m}$ to denote the set of all truthful deterministic mechanisms.
Bundling mechanism. A (deterministic) bundling mechanism sells either all goods together at a bundling price if the sum of the valuations meets or exceeds this price; otherwise, it sells none of the goods. To be precise, it sets a price $p=:\pi(v)$ and defines
$z(v)=\left\{\begin{array}[]{ll}(1,1,\dots,1)&\text{ if }\sum_{i}v_{i}\geq p\\ (0,0,\dots,0)&\text{ if }\sum_{i}v_{i}<p\end{array}\right.$.
2.2 Distributionally robust revenue maximization
In this work, we assume that the valuations of the buyer are random, denoted by the random vector $X=(X_{1},\dots,X_{m})$ over which we have a joint probability distribution $\mathbb{P}^{m}$. We assume that the $X_{i}$ are independently distributed and indicate their marginal distribution by $\mathbb{P}_{i}$, i.e., $\mathbb{P}^{m}$ is the product distribution $\mathbb{P}_{1}\times\mathbb{P}_{2}\times\dots\times\mathbb{P}_{m}$. The (expected) revenue of a mechanism $D$ used for selling $m$ independent goods $(X_{1},\dots,X_{m})\sim\mathbb{P}^{m}$ is given by
$\mathrm{REV}(D,\mathbb{P}^{m})=\mathbb{E}_{\mathbb{P}^{m}}[\pi(X_{1},\dots,X_{m})].$ (2.1)
$\text{BUND}(p_{m},\mathbb{P}^{m})=p_{m}\cdot\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq p_{m}\right).$ (2.2)
In our robust framework, while the joint distribution $\mathbb{P}^{m}$ is unknown, it is assumed to be the product of marginal distributions $\mathbb{P}_{1},\dots,\mathbb{P}_{m}$, each contained in a known common ambiguity set $\mathcal{P}$ of distributions. Although all marginals come from the same ambiguity set, they are not required to be the same. We focus on the mean-MAD ambiguity set (1.1), but describe the revenue maximization problems for the absolute revenue, ratio and regret objective in terms of a general ambiguity set.
The goal is to solve the following maximin and minimax problems. In the maximin problem, the seller first chooses a deterministic mechanism $D\in\mathcal{D}_{m}$, after which an adversary (nature) chooses a distribution $\mathbb{P}^{m}\in\mathcal{P}^{m}$ with the goal of minimizing an objective involving the revenue $\text{REV}(D,\mathbb{P}^{m})$. We consider various objectives, which in this work are sometimes considered for the set of all bundling mechanisms instead of general deterministic mechanisms (but for which the upcoming definitions are the same). In the minimax version of the problem, the setup is the same but this time first nature chooses a distribution, after which the seller chooses a mechanism. We will next give a precise definition of these problems for the objectives of interest.
In the robust (absolute) revenue maximization problem, we want to solve the maximin problem
$\sup_{D\in\mathcal{D}_{m}}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}.$ (2.3)
Intuitively, the goal will be to maximize the average contribution of every good to the revenue. We use this scaling as we will be interested in what happens when $m$ grows large. The minimax problem is given by
$\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\sup_{D\in\mathcal{D}_{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}.$ (2.4)
To define the ratio and absolute regret objective, we need the definition of an optimal deterministic mechanism (i.e., the mechanism that knows the distribution),
$\mathrm{OPT}(m,\mathbb{P}^{m})=\sup_{D^{\prime}\in\mathcal{D}_{m}}\mathrm{REV}(D^{\prime},\mathbb{P}^{m}).$ (2.5)
This optimal mechanism for selling $m$ independent goods with joint distribution $\mathbb{P}^{m}$ is generally not known. Because of individual rationality, we always have $\mathrm{OPT}(m,\mathbb{P}^{m})\leq m\mu$.
In the robust (absolute) regret minimization problem, we want to solve
$\inf_{D\in\mathcal{D}_{m}}\,\sup_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{REV}(D,\mathbb{P}^{m})}{m}.$ (2.6)
That is, we want to bound the absolute difference between the performance of the optimal mechanism that knows the distribution $\mathbb{P}^{m}$ up front, compared to our chosen robust mechanism.
In the robust ratio maximization problem, we want to solve
$\sup_{D\in\mathcal{D}_{m}}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})},$ (2.7)
which is the ratio between the optimal mechanism that knows the distribution $\mathbb{P}^{m}$ and the chosen robust mechanism. This notion is equivalent to the relative regret objective in, e.g., [18]. Note that in (2.7) we do not scale with $m$, because of the multiplicative nature of this objective.
Remark 2.1.
The minimax version
$\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\sup_{D\in\mathcal{D}_{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}$
of the robust ratio maximization problem is trivial, since for any given $\mathbb{P}^{m}$,
$\sup_{D\in\mathcal{D}_{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}=1.$
by definition of $\text{OPT}(m,\mathbb{P}^{m})$ in (2.5). The same argument applies for the regret objective. $\triangle$
In this work, we will often consider a joint distribution of multiple random variables, where each random variable is independently distributed according to a two-point distribution with a given mean $\mu$ and mean absolute deviation (MAD) $d$. The corresponding set of two-point distributions $\mathcal{P}_{2}(\mu,d)\subset\mathcal{P}(\mu,d)$ can be parameterized by the probability mass $\alpha$ on the left support point. The distribution of such a random variable $X(\alpha)\sim\mathbb{P}_{\alpha}$ is given by

### v2

@@ 3.1 Selling below the mean
As our first contribution, Theorem 3.1 shows that, in the many-goods limit, no truthful mechanism can achieve the mean as expected revenue per good. This creates a surprising contrast with the mean-variance ambiguity set, where a straightforward concentration argument for bundling yields the common value of $\mu$.
Theorem 3.1 (Minimax Revenue Theorem).
Let $0\leq d<2\mu$. Then
$\lim_{m\rightarrow\infty}\displaystyle\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{A\in\mathcal{A}_{m}}\,\frac{\mathrm{REV}(A,\mathbb{P}^{m})}{m}\leq\mu-\frac{d}{2}.$ (3.1)
We will next provide a sketch of the main ideas underlying the proof. To analyse the minimax problem in (3.1), we plug in chosen i.i.d. distributions and argue that no (randomized) mechanism can obtain a revenue of more than $\mu-d/2$ as $m$ grows large. To keep the analysis tractable, a first idea is to plug in a two-point distribution $\mathbb{P}_{2}$ supported on $\{x,y\}$ with $\mathbb{P}_{2}(X=x)=\alpha$ and $\mathbb{P}_{2}(X=y)=1-\alpha$. Note that this implies that any realized valuation vector $v$ of the joint distribution over the $m$ goods satisfies $v\in\{x,y\}^{m}$. Because of the mean and MAD constraint that $\mathbb{P}_{2}$ must satisfy, we can fully parameterize two-point distributions in $\mathcal{P}(\mu,d)$ by $\alpha$, i.e., $x=x(\alpha)$ and $y=y(\alpha)$.
Intuitively, the worst-case distribution for the minimax problem corresponds to the situation where $\alpha\rightarrow 1$, which implies that $x(\alpha)\rightarrow\mu-d/2$ (from the left) and $y(\alpha)\rightarrow\infty$. To achieve this, the parameter $\alpha$ is chosen to depend on $m$, i.e., $\alpha=\alpha(m)$, in such a way that $\alpha(m)\rightarrow 1$ as $m\rightarrow\infty$.
By sending $\alpha(m)\rightarrow 1$ at a sufficiently fast speed, the combined contribution of valuation vectors with two or more $y$-valued goods is negligible in the revenue analysis, and so the focus will be on analyzing the contribution of valuation vectors with at most one $y$-valued good. Intuitively, one can argue that the better of setting a deterministic bundling prices just below $m\cdot x(\alpha)$, leading to a revenue of $\mu-d/2$, or a price just below $(m-1)\cdot x(\alpha)+y(\alpha)$, leading to a revenue of $d/2$, is an upper bound on what any mechanism can achieve. In other words, this yields an upper bound of $\max\{\mu-d/2,d/2\}$. For $d<\mu$, so that $\mu-d/2<d/2$, this gives the upper bound of $\mu-d/2$ on the achievable revenue of any mechanism.
To establish the same upper bound for $d>\mu$, it turns out that using a two-point distribution does not suffice. We illustrate this in Appendix A. Instead, we use a more complex distribution that spreads out the probability mass of the point $y(\alpha)$ over $N$ different points $\{x_{1}(\alpha),\dots,x_{N}(\alpha)\}$, with $N$ independent of $m$, in such a way that all points diverge to infinity, but at different speeds as $m\rightarrow\infty$. By doing this carefully the revenue is now, again intuitively, upper bounded by either bundling with a price just below $m\cdot x(\alpha)$ or bundling just below $(m-1)x(\alpha)+x_{i}(\alpha)$ for some $i=1,\dots,N$. Now, however, the revenue in the second case(s) will be at most $d/(2N)$. By choosing $N$ sufficiently large, this means that the first case becomes dominant, thereby establishing the bound of $\mu-d/2$ for $d>\mu$. The full proof is given in Section 4.
To complement Theorem 3.1 we show that deterministic bundling achieves a revenue of $\mu-d/2$ for the maximin problem, thereby establishing that deterministic bundling is optimal in the many-goods limit, and also that the asymptotic maximin and minimax problems have a common value.
Theorem 3.2 (Maximin Bundling Theorem).
Let $0\leq d<2\mu$. Then
@@
$\lim_{m\rightarrow\infty}\displaystyle\sup_{p_{m}\geq 0}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\displaystyle\geq\mu-\frac{d}{2},$ (3.2)
which is asymptotically attained by price
$p_{m}^{*}=(1-\epsilon_{m})^{2}m\left(\mu-\frac{d}{2(1-\epsilon_{m})}\right),$ (3.3)
where $\epsilon_{m}=\frac{1}{2}(1-\frac{d}{2\mu})m^{-1/5}$.
The proof of Theorem 3.2 consists of two main steps. First, we derive a one-sided concentration bound tailored to distributions with finite mean and MAD. Roughly speaking, it states that the sum of $m$ i.i.d. random variables with fixed mean and MAD will be greater than or equal to $\approx m(\mu-d/2)$ with probability close to one as $m\xrightarrow{}\infty$. In the second step, we use this concentration bound to prove the result. The full proof is given in Section 5.
Because bundling attains the upper bound in (3.1), the deterministic maximin and minimax problems coincide in the many-goods limiting setting:
$\lim_{m\rightarrow\infty}\displaystyle\sup_{p_{m}\geq 0}\,\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{REV}(p_{m},\mathbb{P}^{m})}{m}\displaystyle=\lim_{m\rightarrow\infty}\displaystyle\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\displaystyle\sup_{p_{m}\geq 0}\,\frac{\mathrm{REV}(p_{m},\mathbb{P}^{m})}{m}\displaystyle=\mu-d/2.$
It is known that the above relation holds for $m=1$ (Chen et al., 2024, Lemma 1), but this remains unclear when $m>1$, and, in fact, this presents a relevant open problem. We conjecture that equality does not hold, but it remains unclear how to establish this asymptotically. Do notice that, as the bundling mechanism showing optimality for the maximin problem is deterministic, the equality above also holds if one restricts to deterministic mechanisms. In this setting, it is known already for $m=1$ that equality does not hold. This follows from the difference in value between the maximin problem studied in Roos et al. (2021) and the minimax problem in Elmachtoub et al. (2021).
3.2 Price alignment through bundling
Given that deterministic bundling is optimal for maximizing expected revenue in the many-goods limit, we further investigate this subclass of mechanisms for the competitive ratio and the absolute regret objectives. We show that, among all deterministic bundling mechanisms in the many-goods limit, the price that is optimal for expected revenue (3.3) is also optimal for the competitive ratio and absolute regret.
To proceed, we will formalize these objectives. First, the optimal full-information benchmark is
$\mathrm{OPT}(m,\mathbb{P}^{m})=\sup_{A^{\prime}\in\mathcal{A}_{m}}\mathrm{REV}(A^{\prime},\mathbb{P}^{m}).$ (3.4)
The corresponding optimal mechanism is generally not known. Regardless, we do know that $\mathrm{OPT}(m,\mathbb{P}^{m})\leq m\mu$ due to individual rationality. Then, the robust competitive ratio maximization problem is defined as
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})},$ (3.5)
and, analogously, the robust absolute regret minimization problem is
$\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}.$ (3.6)
An interesting open problem is to further investigate the optimality of the bundling price (3.3) within the general class of randomized mechanisms (or important subsets thereof). This, however, will require a different approach than bounding the min-max problem for competitive ratio, since for any $\mathbb{P}^{m}$ (even outside the ambiguity set $\mathcal{P}(\mu,d)^{m}$), we obtain
$\sup_{A\in\mathcal{A}_{m}}\frac{\mathrm{REV}(A,\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}=1$
by definition of $\textup{OPT}(m,\mathbb{P}^{m})$ in (3.4). Consequently, the associated min-max value is always equal to 1. A similar argument can be made for the absolute regret objective.
Theorem 3.3 establishes the universality of deterministic bundling: a single price function is optimal for all three classical objectives in the many-goods limit.
Theorem 3.3 (Unified bundling price).
For $m\xrightarrow{}\infty$, the deterministic bundling mechanism with bundling price
$p_{m}^{*}=(1-\epsilon_{m})^{2}m\left(\mu-\frac{d}{2(1-\epsilon_{m})}\right)$
in (3.3) with $\epsilon_{m}=\frac{1}{2}(1-\frac{d}{2\mu})m^{-1/5}$, asymptotically attains (3.5) and (3.6).
We prove Theorem 3.3 in Section 6 by solving (3.5) and (3.6) separately for $m\xrightarrow{}\infty$. The first gives a guaranteed competitive ratio of $1-d/(2\mu)$, while the second gives an absolute regret of at most $d/2$. We then verify that the bundling price (3.3) is optimal for both objectives. An intuitive way of looking at these results is that the seller can guarantee a revenue close to $m(\mu-d/2)$, whereas the optimal mechanism can achieve a revenue close to $m\mu$, so that the ratio results in $1-d/(2\mu)$ and the absolute regret in $d/2$.
The universality result established in Theorem 3.3 stands in stark contrast to the single-good case. When $m=1$, all three prices for expected revenue, competitive ratio, and absolute regret, denoted by $p^{*}_{rev}$, $p^{*}_{cr}$, and $p^{*}_{ar}$ respectively, are distinct. Theorem 3.4 shows that these prices are strictly ordered, and therefore different. A similar ordering appears in Chen et al. (2022) for the mean-variance ambiguity set.
Proposition 3.4.
Let $0<d<2\mu$. The optimal robust prices for expected revenue, competitive ratio, and absolute regret satisfy
$p^{*}_{rev}<p^{*}_{cr}<p^{*}_{ar}.$ (3.7)
The proof of Proposition 3.4 is found in Section 7. Figure 1 further highlights the contrast with the many-goods setting by showing that $p^{*}_{ar}$ departs markedly from $p^{*}_{rev}$ and $p^{*}_{cr}$. This makes the optimal pricing decision highly sensitive to the choice of objective. In comparison, restricting to deterministic bundling in the many-goods setting resolves this issue.
Figure 1: Optimal robust prices for $m=1$ with $\mu=1$.
4 Proof of the Minimax Revenue Theorem
We now start the proof of Theorem 3.1. We always have for any $A\in\mathcal{A}_{m}$ that
$\lim_{m\rightarrow\infty}\displaystyle\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{REV}(A,\mathbb{P}^{m})}{m}\leq\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{A\in\mathcal{A}_{m}}\,\frac{\mathrm{REV}(A,\mathbb{P}^{m})}{m}\leq\lim_{m\rightarrow\infty}\sup_{A\in\mathcal{A}_{m}}\,\frac{\mathrm{REV}(A,\mathbb{P}^{m}_{N})}{m},$
where we choose $\mathbb{P}^{m}_{N}=\Pi_{i=1}^{m}\mathbb{P}_{N}$ with $\mathbb{P}_{N}$ as